OpenStudy (anonymous):

pythagorean theorem c= 13 b=9 a=?

OpenStudy (anonymous):

9.38

OpenStudy (anonymous):

a^2=b^2+c^2 a^2=169+81=250 a=5root(10)

OpenStudy (anonymous):

5\[\times\]\[\sqrt{10}\]=15.811 dhashni is correct...

OpenStudy (anonymous):

Dhashini and Electronic Rajesh how can u assume the sides like that.....it not given where the right angle is

OpenStudy (anonymous):

kyra will be better if u give a diagram for the same the next time

OpenStudy (anonymous):

ok jetly if you assume c to be hypotenuse .. then your answer is right... but if a is hypotenuse then dhashni is correct..

OpenStudy (anonymous):

that is why i asked kyra for a diagram the next time......sorry but no offence

OpenStudy (anonymous):

don't feel like that...